OpenStudy (anonymous):

Please help! "State the period of the following functions" - g(x) = 4sin(pie x+2)

OpenStudy (badhi):

Consider a general sinusoidal function = \[g(x) = \sin(\omega x+\phi) \] the period of the wave will be = \(\frac{2\pi}{\omega}\)